OpenStudy (anonymous):

[4.06] What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

Graph it @ https://www.desmos.com/calculator

OpenStudy (anonymous):

Copy & Paste all 4 equations in seperate boxes on the left. Then see which one crosses both points (-2, 1) and (1, 10).

OpenStudy (anonymous):

*separate

OpenStudy (anonymous):

I've already found the answer.

OpenStudy (texaschic101):

first we find the slope slope(m) = (y2 - y1) / (x2 - x1) (-2,1) x1 = -2 and y1 = 1 (1,10) x2 = 1 and y2 = 10 now we sub slope(m) = (10 - 1) / (1 - (-2) slope(m) = 9/3 slope(m) = 3 now we use y = mx + b slope(m) = 3 you can use either of your sets of points (1,10) x = 1 and y = 10 now we sub 10 = 3(1) + b 10 = 3 + b 10 - 3 = b 7 = b our equation is : y = 3x + 7 but we need it in standard form Ax + By = C y = 3x + 7 -- subtract 3x from both sides -3x + y = 7 -- multiply by -1 to make x positive 3x - y = -7

OpenStudy (texaschic101):

any questions at all ?

OpenStudy (anonymous):

thank you texaschic, can you help me on one more

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

In y = mx + b form, the m is the slope and the b is the y intercept 3x + 2y = 12 ---- subtract 3x from both sides 3x - 3x + 2y = -3x + 12 2y = -3x + 12 -- divide both sides by 2 y = -3/2x + 6 slope(m) = -3/2, and y intercept (b) = 6 or (0,6)

OpenStudy (anonymous):

ok, which answer is it?

OpenStudy (texaschic101):

first one

OpenStudy (anonymous):

thanksssssss

OpenStudy (texaschic101):

sure thing :)